OpenStudy (anonymous):

If there are 1200 calories in 8 ounces of hot fudge, how many calories are in 3 ounces of hot fudge

OpenStudy (anonymous):

450, you have to divide the amount 8 ounces has and find out how much 1 has and multiply it by 3

OpenStudy (anonymous):

1200/8 = 150 thus, 150 per calorie. 150 x 3 = 450 :)

OpenStudy (anonymous):

At a certain college the ratio to men to women is 6 to 5. If there are 1500 men, how many women are there

OpenStudy (anonymous):

6 / 5 = 1500 / X then cross multiply X = 5 * 1500 / 6

OpenStudy (anonymous):

philips13 I got 450 thank you!(;

OpenStudy (anonymous):

that's the fudge, I was responding to the college men to women ratio

OpenStudy (anonymous):

Ohh

OpenStudy (anonymous):

\[{6 \over 5 }={ 15000 \over x }\]